OpenStudy (anonymous):

Given the function f(x) = log3(x + 4), find the value of f−1(3). f−1(3) = 7 f−1(3) = 12 f−1(3) = 23 f−1(3) = 31

OpenStudy (loser66):

let log 3(x+4) =3 to find x

OpenStudy (loser66):

\(f(x) =\color{red}{a}\), then \(f^{-1}\color{red}{(a)}= x\) hence if you let \(log_3(x+4)=\color{red}{3}\) then \(f^{-1}\color{red}{(3)}=x\) so, just solve for x

OpenStudy (anonymous):

so we make (x+4)^log3=3 and we subtract 4 from both sides? @Loser66

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@i_suck_at_school

OpenStudy (anonymous):

@quickstudent

OpenStudy (anonymous):

I justHAD that question

OpenStudy (anonymous):

hang on\

OpenStudy (anonymous):

are you taking alg2 ?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

can you help me solving this? please

OpenStudy (quickstudent):

I'm not at 2 yet. I'm still studying algebra 1.

OpenStudy (anonymous):

@Preetha @mathrulezz @hartnn

OpenStudy (anonymous):

is it a test

OpenStudy (anonymous):

check this out and see if it helps

OpenStudy (anonymous):

http://www.math.utah.edu/~wortman/1050-text-l.pdf

OpenStudy (anonymous):

i appreciate it @i_suck_at_school but idk it didn't answer my question , i'm so frustrated

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

@Luigi0210 @Directrix @sammixboo @Miracrown @Legends @OpTicRebekah @Data_LG2 @Shadowgirl15 @IrishBoy123 @kourtneyceciliax

OpenStudy (anonymous):

f (x) = log3 (x −2) find f (11) y = log3(11− 2) y = log3 (9) to find log3 [ (9) ask what power of 3 is 9 log3(9) = 2] y = log3(9) y = 2 (11 ,2 )

OpenStudy (anonymous):

not you problem just example

OpenStudy (irishboy123):

i assume that f-1(3) means the inverse. this has been presented badly and may have led to confusion. IF it is true that you are trying to find an inverse, then you are trying to solve: log3(x + 4) = 3, ie what x gives 3 as the answer assuming base e (as opposed to 10 or any other poss base), this means that: 3(x+4) = e^(3) solve from here for x. assuming base 10, you want to solve: 3(x+4) = 10^3 hope this helps.

OpenStudy (anonymous):

it is c i am sorry 4 giving answer

OpenStudy (anonymous):

too late but here's how i do it: let f(x)=y solve for the inverse... remember inverse is switching x and y \(\sf f(x) = log_3(x + 4)\\y=log_3(x + 4)\\\color{red}{x}=log_3(\color{red}{y} + 4)\) Arrange the equation in terms of \(\sf \color{red}{y}\) then plug the give value of x, which is 3 you'll end up the same thing as they got :)

OpenStudy (anonymous):

@DoShKa_SyRiA

OpenStudy (anonymous):

Thank you so much i understand it now @Data_LG2

OpenStudy (anonymous):

@i_suck_at_school i appreciate your help a lot man

OpenStudy (anonymous):

@i_suck_at_school how is it c i thought it was A.

OpenStudy (anonymous):

i am a girl

OpenStudy (anonymous):

oh i'm sorry you have a guy picture so that's why i assumed so

OpenStudy (anonymous):

@i_suck_at_school

OpenStudy (anonymous):

its ight

OpenStudy (anonymous):

but why did you say it's c?

OpenStudy (anonymous):

isn't the answer A?

OpenStudy (anonymous):

so we plug in 3 in x .. 3= log3(y+4) correct?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

3 = log3(x + 4) x + 4 = 3^3 = 27 ... i don't understand this part x = 23 f-1(3) = 23.

OpenStudy (anonymous):

@i_suck_at_school @Data_LG2

OpenStudy (anonymous):

http://www.purplemath.com/modules/logrules3.htm

OpenStudy (anonymous):

scroll to bottom and ask away

OpenStudy (anonymous):

@DoShKa_SyRiA

OpenStudy (anonymous):

g2g c is right

OpenStudy (anonymous):

just sayin